The weights of a women population are normally distributed with mean 60kg and standard deviation 5kg. Find the probability that the weight of a woman chosen at random is more than

65kg.


1
Expert's answer
2021-05-07T12:38:22-0400

To determine the probability of making options from random choices, the following expression is utilized: "X-N\left(\mu ,\:\sigma ^2\right)\:hence\:Z=\frac{X-u}{\sigma }-N\left(0,\:1\right)" provided that weight"W-N\left(60,\:5^2\right)" then "Z=\frac{W-60}{5}-N\left(0,\:1\right)\:"


To calculate "P\left(W>65\right)"

"P\left(W>65\right)\:=P\left(\frac{W-60}{5}>\frac{65-60}{5}\right)"

"=P\left(Z>1\right)"

"=1-0.9772"

"=0.0228\:"